# SonarSplat: Novel View Synthesis of Imaging Sonar via Gaussian Splatting

### v1

@@ 4.4 Azimuth Streak Removal
We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a pointcloud for visualization. We compare with methods focusing on 3D reconstruction: Neusis and DSC [30, 7]. Qualitative results are shown in Fig. 7 and quantitative results are shown in Tab. 4. We use popular metrics from scene reconstruction literature and choose $l1$ Chamfer Distance, F-Score, and voxelized IoU to evaluate performance [1]. For Neusis and DSC, we produce meshes using marching cubes. For SonarSplat, we choose to sample points from the 3D Gaussians to produce a pointcloud. We sample 100,000 points from both the mesh and Gaussian splat and compute evaluation metrics against a ground truth photogrammetry reconstruction of the rock platform. We note that neither Neusis nor DSC are able to reconstruct the object whereas SonarSplat is able to produce a coarse representation.
@@
Method CD-$l_{1}$ (m) $\downarrow$ F-Score $\uparrow$ IoU $\uparrow$
Neusis [30] 0.202 0.4834 0.0129
DSC [7] 0.187 0.4613 0.0164
SonarSplat (ours) 0.084 0.1170 0.0191
Table 4: We quantitatively evaluate the 3D reconstruction of selected methods on the rock_360 sequence. Best is in bold.
5 Conclusion & Future Work
We propose SonarSplat, the first sonar-only Gaussian splatting framework for novel view synthesis for imaging sonar in underwater applications. First, we adapt the sonar rendering equation for efficient range-azimuth splatting by evaluating 3D Gaussians. We model azimuth streaking within the Gaussian splatting framework, allowing us to learn per-Gaussian azimuth streaking probabilities and produce high-quality de-streaked images. We perform experiments on real-world data in test tank and river environments. In the novel view synthesis task, SonarSplat outperforms baselines by +2.5 dB PSNR and demonstrates superior qualitative results. We also demonstrate that SonarSplat learns scene geometry, demonstrated by both quantitative and qualitative results. Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability. This can yield more diverse datasets for training sonar-based scene understanding algorithms. We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise, Lambertian scattering, and multi-path reflections.
6 Acknowledgment
This work is supported by the National Science Foundation under Award No. 2337774.

### v2

@@ IV-F Real-World 3D Reconstruction
We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a mesh for visualization. We compare to state-of-the-art sonar reconstruction methods: Neusis, DSC, and ZSplat* [6, 33, 35]. Qualitative results are shown in Fig. 8 and quantitative results are shown in Tab. IV. Following [33, 6], we report Root Mean Squared (RMS) $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). We produce meshes from all methods then randomly sample 30,000 points from the surface for 30 iterations to calculate RMS. Note that neither DSC, Neusis, nor ZSplat* is able to reconstruct the object whereas SonarSplat is able to produce an accurate reconstruction of the piling.
TABLE IV: Quantitative evaluation of 3D reconstruction performance on the piling using RMS $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). Best in bold.
@@
Method CD-$l_{1}$downarrow HD $\downarrow$
Neusis [33] 0.32 0.78
DSC [6] 0.29 0.82
ZSplat* [35] 2.93 4.88
SonarSplat (ours) 0.14 0.57
V Conclusion & Future Work
We propose SonarSplat, the first sonar-only Gaussian splatting framework for imaging sonar in underwater applications. First, we adapt the sonar rendering equation for efficient range/azimuth splatting by evaluating 3D Gaussians. We model azimuth streaking within the Gaussian splatting framework, allowing us to learn per-Gaussian azimuth streaking probabilities and produce high-quality de-streaked images. We also introduce a novel densification strategy (ESDS) to enable SonarSplat to resolve elevation angle ambiguities and associated occlusions. From our extensive experiments on a diverse set of real-world data in test tank and river environments, we find that SonarSplat outperforms state-of-the-art methods in novel view synthesis (+3.2 dB PSNR) and 3D reconstruction (52% lower Chamfer Distance), and additionally allows for de-streaking of sonar images.
Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability. This can yield more diverse datasets for training sonar-based scene understanding algorithms, including multi-view target recognition algorithms [40, 38]. We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise, Lambertian scattering, and multi-path reflections. Our choice of Gaussian splatting as a 3D scene representation also opens possibilities for real-time sensor fusion and SLAM [13, 23].